Find the formula dor the inverse of the given function, if possibleW(x)=(x^3+4)^1/5-1

Answer by htmentor(1343)

Find the inverse of the function W(x)=(x^3+4)^1/5-1
==============
W(x) = y = (x^3+4)^1/5-1
To find the inverse, interchange x and y:
x = (y^3+4)^1/5 - 1
Solve for y:
(y^3+4)^1/5 = x+1
y^3 + 4 = (x+1)^5
y^3 = (x+1)^5 - 4
y = f_inv(x) = [(x+1)^5 - 4]^(1/3)